### v2

# Adaptive Data Collection for Latin-American Community-sourced Evaluation of Stereotypes (LACES)

Abstract
The evaluation of societal biases in NLP models is critically hindered by a geo-cultural gap, This leaves regions such as Latin America severely underserved, making it impossible to adequately assess or mitigate the perpetuation of harmful regional stereotypes in language technologies.
This paper presents LACES, a stereotype association dataset, for 15 Latin American countries. This dataset includes 4,789 stereotype associations manually created and annotated by 83 participants. The dataset was developed through targeted community partnerships across Latin America.
Additionally, in this paper, we propose a novel adaptive data collection methodology that uniquely integrates the sourcing of new stereotype entries and the validation of existing data within a single, unified workflow. This approach results in a resource with more unique stereotypes than previous static collection methods, enabling a more efficient stereotype collection. The paper further supports the quality of LACES by demonstrating reduced efficacy of debiasing methods on this dataset in comparison to existing popular stereotype benchmarks.
Adaptive Data Collection for Latin-American Community-sourced Evaluation of Stereotypes (LACES)
Guido Ivetta1,2, Pietro Palombini1, Sofía Martinelli1, Marcos J. Gomez1,2, M. Emilia Echeveste1,2, Sunipa Dev3, Vinodkumar Prabhakaran3, Luciana Benotti1,2 1Universidad Nacional de Córdoba, Argentina 2Fundación Vía Libre 3Google Research guidoivetta@unc.edu.ar
1 Introduction
While mitigating the social biases that NLP models learn from web-scale training data continues to remain a challenge, the resources that we rely on to evaluate these harms are themselves critically biased. A growing body of research demonstrates that NLP models tend to perpetuate and sometimes amplify social stereotypes Bolukbasi et al. (2016); Garg et al. (2018), underscoring the need for robust stereotype evaluation datasets; yet, the field is dominated by resources that are overwhelmingly English-centric and focused on U.S. demographics. While NLP researchers have brought attention to these gaps in recent years Prabhakaran et al. (2022), glaring gaps continue to persist in the global coverage of evaluation resources, especially in Latin America.
Figure 1: This dataset covers 4789 stereotypes, covering 120 identities and 842 attributes. It was built by 83 annotators from 15 distinct countries. The map illustrates the participating Latin American nations. The methodology of collection adapts to the participant identity by bringing examples relevant to their nation of origin.
Recent research has begun to address these gaps in evaluation data by curating candidate stereotypes from large language models themselves Bhutani et al. (2024); Jha et al. (2023), although this approach anchors primarily on the data and sociocultural knowledge captured in the training set and fails to include more nuanced and broader community insights. In contrast, there is also recent complementary work that engages with a larger set of participants (e.g., Dev et al. (2023); Maina et al. (2024); Mitchell et al. (2025); Ivetta et al. (2025)) to collect a broader set of stereotypes. Our work is in line with these efforts, with a focus on sourcing stereotypes across countries in Latin America.
These community-sourced data collection approaches tend to be static and non-adaptive — i.e., they are agnostic to the existing state of knowledge, offering no mechanism to dynamically assess their own comprehensiveness or coverage. Consequently, when sampling data from a specific locale, we cannot systematically identify critical gaps or data sparsity. This leaves datasets often containing redundant information about the most common cultural knowledge, while the long-tail information is often missed. The resulting datasets will be incomplete in unknown ways and provides no guarantee that we have successfully captured the necessary spectrum of cultural knowledge.
This “blind sampling” methodology creates a second, concurrent failure: profound resource inefficiency due to data redundancy. Lacking a mechanism to track known information, collection efforts repeatedly oversample the high-frequency “head” of the cultural knowledge distribution—typically the most salient, common-knowledge concepts. It not only wastes resources by re-annotating the known “head” but also fails to capture the diverse, nuanced knowledge in the “long tail.” Critically, this ad-hoc redundancy is currently treated as a mere annotation artifact and discarded, rather than as a rich sociolinguistic signal. Systematically captured repetition, for instance, serves as a powerful proxy for saliency (or prevalence) and pluralism (contested perspectives).
In this paper, we describe a novel adaptive data collection approach that we utilized for collecting stereotype data across 15 countries in Latin America, collected over two separate initiatives. We demonstrate how our adaptive approach results in reduced redundancy, while improving coverage above and beyond what a traditional data collection approach would have yielded. Our resulting dataset has 4789 stereotypes, covering 120 identities and 842 attributes across a multitude of different sociodemographic axes, as annotated by 83 individuals from 15 distinct countries. We also demonstrate critical gaps in stereotype evaluations that our dataset reveals, and that the existing mitigation approaches crucially overlooks these stereotypes.
2 Related Work
In this section we first review previous work on stereotype dataset creation and explore how datasets are not neutral repositories of linguistic knowledge but are shaped by the choices of stakeholders involved. Then we review how examples can shape the data collected particularly in the creation of stereotype datasets. Finally we describe methods that have been used to dynamically adapt the examples so that they are meaningful and relevant for the annotators.
Existing stereotype datasets
Several benchmark datasets have been proposed to quantify and analyze stereotypes encoded in language models. Early efforts such as CrowS-Pairs (Nangia et al., 2020) and StereoSet (Nadeem et al., 2021) focused on English and U.S.-centric contexts, measuring stereotypical bias across domains like race, religion, gender, and profession. BBQ (Parrish et al., 2022) extended this line of work to question answering, showing that models tend to rely on stereotypes when contextual information is under-specified. More recent resources broaden geographic and linguistic scope. SeeGULL (Jha et al., 2023) used language models to propose stereotypes from different regions and validated them with globally diverse annotators to construct a dataset covering stereotypes from 178 countries across six continents, while SHADES (Mitchell et al., 2025) provided a multilingual parallel dataset spanning 16 languages and 20 regions. Complementary work has emphasized participatory and community-driven approaches: Dev et al. Dev et al. (2023) engaged with Indian communities to surface locally grounded stereotypes absent from Western-centric benchmarks. Ivetta et al. Ivetta et al. (2025) is a similar effort that engaged Latin American communities. These newer datasets highlight the need for socio-culturally inclusive stereotype evaluation. However, all of them use fixed examples to elicit stereotypes instead of adapting the examples to the social group of the annotator, as we do in this paper.
Datasets as shaped artifacts
Different annotators will not necessarily assign the same labels to the same texts, resulting in human label variation Plank (2022a). There is evidence that this variation depends on the demographic characteristics of annotators (Binns et al., 2017; Al Kuwatly et al., 2020; Excell and Al Moubayed, 2021; Shen and Rose, 2021). Variation is stronger for subjective tasks like toxic content detection (Sap et al., 2019; Kumar et al., 2021; Sap et al., 2022; Goyal et al., 2022) and stereotype elicitation. Annotation guidelines are known to influence the resulting data obtained. In particular, for concepts such as safety and offensiveness Davani et al. (2024); Aroyo et al. (2023) the definitions of the annotation task are subjective and can be interpreted in different ways. Different definitions of bias and stereotypes have been discussed in previous work, which has found that frequently the definitions are inconsistent or inexistent in work related to these topics in NLP. These definitions are routinely accompanied by examples that illustrate the definition on which annotators rely to understand the definition and from which they generalize. Annotation guidelines play a particularly influential role: the examples included in these documents often serve as prototypes that shape annotators’ interpretations of the task Rogers (2021). Previous work has called for the documentation of data collection guidelines including its examples Bender and Friedman (2018).
Examples as a source of bias
Another underexplored but critical issue concerns how annotation guidelines themselves shape the form of datasets. Prior work in data documentation and dataset design highlights that guidelines choices, including the selection of illustrative examples, are not neutral but actively steer annotators towards certain interpretations and away from others Gebru et al. (2021); Paullada et al. (2021). It has been argued that these seed examples Antoniak and Mimno (2021) are a brittle but unavoidable element of current data collection, particularly when no clear crystal definitions and ontologies are complete. When examples are pre-selected by researchers, they risk embedding the researchers’ own cultural assumptions and biases into the dataset. In contrast, dynamically adapting examples based on the contributions of the annotators’ own social groups can reduce top-down bias and surface more authentic, community-grounded knowledge. Such an approach resonates with scholarship on participatory methods in dataset creation Miceli et al. (2022); Jo and Gebru (2020); Zhao et al. (2024), which argues for shifting epistemic authority from researchers to annotators and their communities.
Adaptive stereotype collection
Rather than treating disagreement among annotators as noise, recent work has argued for leveraging such variation to capture multiple perspectives Davani et al. (2022). Complementary to these efforts, research on dynamic example presentation has explored how to support annotators with more contextually relevant references. For instance, the DEXA framework Zlabinger et al. (2020) proposes providing annotators with semantically similar examples drawn from expert annotations during the labeling process. This approach reduces reliance on static guideline examples and points towards a more adaptive model of annotation support, though it has not yet been extended to capture the social dynamics of groups contributing their own examples. Existing research has treated redundancy in annotations as noise to be discarded, but as argued in Aroyo and Welty (2015), repeated signals can in fact serve as valuable indicators of saliency, prevalence, or contested interpretations. By explicitly capturing and adapting to redundancy, our approach reduces inefficiency while improving both breadth and depth of stereotype coverage.
3 Data Collection Methodology
In this section, we first describe the physical and virtual contexts in which the data was gathered. Then, we detail the data collection task, interactive interface, and adaptive sampling strategies employed to build the dataset.
3.1 Data Collection Contexts
The data collection was conducted in two settings. First, we organized an in-person workshop at Khipu 2025 (https://khipu.ai/). We adopted a simultaneous and co-located strategy, where all participants interacted with the tool in real time. Unlike conventional crowd-working, participants were physically present in the same room, performing the task simultaneously.
The group included both members of the research community with prior NLP training and newcomers to the field. The workshop lasted 2 hours, facilitated by 3 people with educational training in NLP and ethics. To ensure anonymity while maintaining traceability, each participant was assigned a random identifier not linked to any personal information.
Second, the task was also integrated as an activity within an SomosNLP Hackathon (https://somosnlp.org/), extending the data collection process to an online environment. This allowed us to reach a larger pool of contributors across Latin America and Spain, complementing the in-person workshop, and providing additional validation and new associations.
3.2 Data Collection Approach
The core task for data collection consisted of validating a given (nationality, attribute) pair, and extending the dataset by introducing new pairs, either by proposing a new attribute for the nationality or by associating another nationality to the attribute.
The annotation interface is illustrated in Figure 2. At the top, participants were presented with a sample pair of (nationality, attribute), where the nationality appeared in red and the attribute in green. Immediately below, they were asked to evaluate the statement “This is a known association in my region” using a 5-point Likert scale. Following this, participants could provide additional pairs through the optional fields shown in the interface. In (Figure 2, Brazil), they could select other nationalities that they believed were also stereotypically associated to the given attribute. In (Figure 2, make strangers feel like family), participants could propose new stereotypical attributes for the nationality displayed in the initial pair.
Our methodology is characterized by its adaptive nature. Each new pair generated by participants was automatically added to the pool of items available for others to evaluate. This mechanism created a continuous feedback loop in which participants not only validated existing associations but also expanded the dataset by observing what other participants added previously.
Before the annotation phase, we collected participant demographics, including country of origin, cultural affiliations, and language proficiency. The selection of input pairs for each participant was governed by a weighted probabilistic sampling algorithm rather than a hard constraint. This approach was designed to balance three primary objectives:
1. In-group Representation: The algorithm weighted geographically or culturally proximate pairs more heavily, ensuring participants primarily evaluated data from their own communities.
2. Validation Coverage: Priority was given to pairs with fewer than three existing validations.
3. Setting-specific Recency: The system prioritized data from the current session to foster real-time feedback and peer engagement.
By employing a probabilistic framework instead of rigid rules, the system maximized in-group validation while maintaining a balanced distribution of annotations, thereby enhancing the overall depth and robustness of the dataset. To initialize the task, we relied on a small and manually curated seed set derived from the HESEIA dataset Ivetta et al. (2025) which focuses in Argentina.
Figure 2: Interface of the data collection tool showing the (nationality, attribute) pair, Likert scale validation, and optional fields for additional datapoint associations.
The system only presented pairs in the language the annotator reported understanding, and annotations could be done in any of their declared languages. This setup enabled the construction of a multilingual dataset. Although Spanish dominated due to the demographics of the participant pool, the design allowed for the integration of examples in multiple languages. Furthermore, participants could decide not to annotate a pair by skipping it.
Topic n IG % IG association sample OG % OG association sample
Cooking and Food 792 39.64 (CHL, piscola) 60.36 (PRY, tortafrita)
Positive Traits 641 27.78 (URY, hospitable) 72.22 (JPN, problem solvers)
Geography, Buildings, Landmarks 609 26.21 (MEX, archaeology) 73.79 (BRA, Cristo Luz)
Economy 591 8.88 (PER, cheap tourism) 91.12 (AUS, work & holiday)
People & Everyday Life 571 13.08 (CRI, ecological) 86.92 (CHN, work culture)
Tradition, Art, History 388 26.27 (CHL, rodeo) 73.73 (GRC, sirtaki)
Negative Traits 338 23.66 (COL, fallacious) 76.34 (DEU, rigid minded)
Politics & Governance 239 18.57 (ARG, best education) 81.43 (CUB, public health)
Sports & Recreation 223 50.65 (COL, football fans) 49.35 (RUS, athletes)
Other 137 16.67 (BRA, dental health) 83.33 (ISL, attractive people)
Public Figures & Pop Culture 130 63.64 (URY, Gardel) 36.36 (CUB, Fidel Castro)
Neutral Traits 130 34.57 (PAN, serious) 65.43 (IRL, quiet people)
Table 1: Topics in the LACES dataset. For each topic, the table shows the total frequency ($n$), the percentage of associations created by annotators from in-group (IG) and out-group (OG) perspectives, and a representative example of each. Countries are represented with their ISO 3-letter code.
While our implementation focused on language and nationality, the same dynamic mechanism could be extended to other participant characteristics or research objectives, enabling diverse data collection strategies. To facilitate reproducibility and support future research, we have released the source code as an open-source framework. The architecture allows researchers to modify the adaptive sampling logic and the annotation interface to their own data collection needs, providing a versatile infrastructure for various types of community-based, human-in-the-loop initiatives.
4 Dataset Characterization
This dataset covers 120 identities and 842 attributes as annotated by 83 individuals from 15 distinct countries (see Appendix A for nationality distribution). Annotators often possessed multiple language proficiencies: 92% claimed to read and write English, apart from Spanish or Portuguese. The dataset is multilingual and comprises 2437 (nationality, attribute) pairs in Spanish, and 2352 in English.
We employed an adaptive validation methodology, resulting in a dynamic distribution of ratings across the dataset. While all entries attained a minimum of one validation, a subset of 765 unique pairs in Spanish and 501 in English were validated at least twice. This approach allows for real-time sampling and resource optimization, enabling us to prioritize data points based on specific research objectives: incentivizing geographic proximity, capturing under-explored topics, ensuring diverse demographic representation, or increasing the validations for controversial topics to better quantify agreement.
In this section we will first present the explored topics in LACES with supporting examples. Then, we examine the data validation and resulting inter-annotator agreement, determining which topics demonstrated the strongest consensus and the greatest disagreement. Finally, we analyze in-group bias in annotator ratings, showing that people agree more when rating positive and neutral attributes associated with their own group.
4.1 Topics Explored in LACES
To identify the main topics represented in our dataset, a categorization scheme was devised, inspired by the CVQA Romero et al. (2025) and OK-VQA Schwenk et al. (2022) benchmarks. Six additional categories were introduced to better fit the scope of the project: Politics & Governance, Economy, Positive Traits, Negative Traits, Neutral Traits, and Other. Category assignment was manually reviewed and refined.
Table 1 shows the set of categories, their frequency, illustrative examples, and the percentage of associations created from in-group (IG) and out-group (OG) perspectives; reflecting whether annotators referred more to their own group or to others. Given that 15 countries participated in the study, a purely random sampling approach would yield only 6.67% coverage for in-group annotations. However, since our adaptive methodology specifically incentivized their collection, IG percentages significantly exceed this baseline across all categories.
4.2 Inter-Annotator Agreement
To identify which categories elicit higher levels of disagreement, and may thus be considered more controversial, we quantified inter-annotator agreement by calculating the variance of scores assigned to each unique (nationality, attribute) pair. High variance in these ratings serves as a proxy for sociocultural contention within a given topic.
For this analysis, only pairs with at least two validations were retained. Approximately 40% of pairs had zero variance, and the 75th percentile corresponded to a variance below 1. Pairs were classified into two groups: low variance (80%) and high variance (20%). We did not adopt a finer-grained partition since smaller groups would have contained very few samples, complicating the analysis and potentially adding noise into the interpretation.
We analyzed whether annotator disagreement was associated with specific topics. Relative change for each category between the two groups was computed as $100\times\frac{\text{High}-\text{Low}}{\text{Low}}$, quantifying how much more frequent each category was in the high-variance group.
Figure 3 depicts these relative frequency changes. A large positive change indicates high controversy, whereas a negative change suggests consensus. Notably, Negative Traits emerged as the most controversial category, with a relative change of 182.98%, meaning it was almost three times more prevalent in the high-variance group than in the low-variance group. By contrast, positive and neutral traits did not exhibit comparable patterns of disagreement. On the other side, Sports & Recreation shows the most consensus across topic categories.
Figure 3: Relative change from low to high-variance group across topics. A positive relative change indicates that a category becomes more frequent in the high-variance group (controversial), while a negative relative change indicates that it is more frequent in the low-variance group (consensus).
4.3 In-Group Bias and Self-Attribution
To further understand how social identity influences the perception of stereotypes, we manually augmented the attributes in the LACES dataset with sentiment labels: Positive, Neutral, or Negative. This allows us to analyze the relationship between an annotator’s group membership and their willingness to validate a stereotype. Figure 4 illustrates the distribution of responses to the prompt: “This is a known association in my region” measured on a 5-point Likert scale (1 = completely unknown, 5 = very well-known).
Figure 4: Distribution of average stereotype recognition scores (1 = unknown, 5 = very well-known) for ratings of an annotator’s own group (blue) versus other groups (orange), across attribute sentiment. Annotators report a higher prevalence of positive and neutral attributes for their own community.
The results reveal an in-group leniency effect mediated by the sentiment of the attribute. When participants evaluated stereotypes targeting their own nationality (In-Group), we observed that Positive and Neutral attributes showed a distinct shift toward the top of the scale, with mean recognition scores of $4.27$ and $4.20$, respectively. In contrast, participants were notably less likely to acknowledge the social prevalence of Negative stereotypes about their own group, with the mean score dropping to $3.51$. This suggests that individuals are more willing to validate the social prevalence of favorable or benign traits within their own community, likely reflecting a form of collective defensive self-attribution or social desirability bias.
In contrast, the perception of out-group stereotypes showed almost no variation across the three sentiment categories, with means of $3.53$ for Positive, $3.57$ for Neutral, and $3.43$ for Negative attributes. This indicates that while in-group stereotypes are deeply partitioned by the desire to maintain a positive social identity, out-group stereotypes are viewed through a more detached and generalized lens.
By capturing this divergence, LACES offers a unique perspective on the intersection of sentiment and identity in data annotation, highlighting that even well-known stereotypes are filtered through the lens of one’s own community membership.
5 Analysis and Benchmark Comparison
This section compares our dataset with existing benchmarks. First, we quantify the dataset conceptual diversity, indicating it contributes a high percentage of unique concepts not covered by similar resources. Then, we evaluate the performance of self-debiasing methods.
5.1 Unique Attributes
To quantify the conceptual novelty of LACES, we measure its thematic overlap with stablished and recent benchmarks: CrowS-Pairs (Nangia et al., 2020), BBQ (Parrish et al., 2022), SeeGULL (Jha et al., 2023), SHADES (Mitchell et al., 2025), and HESEIA (Ivetta et al., 2025).
Our methodology employs a semantic vector space approach to identify thematic redundancies rather than relying on exact keyword matching. We transformed each data point into a high-dimensional representation using the multilingual text-embedding-3-large model (OpenAI, 2026). For every entry across all datasets, we calculated the maximum cosine similarity against all entries in the comparison pool. An attribute is classified as "unique" if its embedding does not exceed a predefined similarity threshold with any existing record in the other datasets, indicating that the entry represents a novel thematic contribution.
As shown in Table 2, the results demonstrate the percentage of unique concepts within each dataset when analyzed for thematic overlap, revealing that the LACES dataset contains the highest percentage of unique concepts at 29.74%. This figure represents an improvement over other regional and global benchmarks, surpassing HESEIA (27.78%) and SeeGULL (19.63%), and significantly over the uniqueness found in other resources such as BBQ (15.35%), CrowS-Pairs (14.33%), and SHADES (10.85%). The overall conceptual overlap between datasets can be visualized in Appendix C.
Dataset Size Unique Attributes
LACES 4,879 1,650 29.74%
HESEIA 45,416 13,134 27.78%
SeeGULL 25,861 6,716 19.63%
BBQ 58,492 11,452 15.35%
CrowS-Pairs 1,508 321 14.33%
SHADES 728 79 10.85%
Table 2: Comparison of Dataset Size and Attribute Diversity. This table presents size, total unique attributes, and the percentage of unique concepts (highest bolded) across evaluated datasets. The uniqueness percentage is calculated using thematic similarity embeddings.
Interestingly, while both participatory benchmarks (HESEIA and LACES) contain the highest proportions of unique concepts, LACES proves more efficient. HESEIA took approximately 8,535 hours to collect (1.5 hours in average per 5,690 participants) while LACES took approximately 300 hours (3.6 hours in average per 83 participants). Despite LACES containing ten times fewer data points than HESEIA and only a small fraction of participants, it yields a higher percentage of unique entries. We hypothesize that this is a result of our adaptive methodology, which actively incentivizes annotators to explore the "long tail" of sociocultural knowledge. In contrast, the static examples provided in HESEIA and other resources, in the annotation guidelines, may have inadvertently constrained annotator exploration to more salient attributes.
5.2 Self-Debiasing
To evaluate the generalizability of zero-shot self-debiasing, we extend techniques proven effective in BBQ to the datasets analyzed in the previous section. HESEIA was excluded as it does not distinguish between stereo- and anti-stereotypes, which is required for this experiment. We adopt the BBQ evaluation protocol, using multiple-choice items that ask which country is associated with a specific stereotypical attribute, including an "Unknown" option to mitigate forced-choice bias. Following Gallegos et al. (2025), we evaluate three debias strategies (see Appendix B for details):
Baseline (B): A direct prompt for a single-letter answer without reasoning or intervention.
Explanation (E): An approach where the model identifies invalid assumptions before answering.
Reprompting (R): A two-step intervention instructing to remove bias from its initial response.
Parrish et al. (2022) defines $\textsc{Bias}=(1-\textsc{ACC})\left[2\frac{n_{\textsc{biased}}}{m}-1\right]$, ACC is accuracy, $m$ is the number of non-Unknown predictions, and $n_{\textsc{biased}}$ counts predictions matching the bias target.
Dataset S GPT Llama-3.1
3.5-Turbo 4o-mini
LACES B 0.53 0.68 0.55
E 0.23 0.52 0.38
R 0.34 0.20 0.39
SeeGULL B 0.27 0.31 0.24
E 0.09 0.15 0.13
R 0.12 0.02 0.11
BBQ B 0.12 0.14 0.10
E 0.03 0.01 0.00
R 0.04 0.01 0.02
CrowS-Pairs B 0.18 0.15 0.18
E 0.04 0.01 0.03
R 0.09 0.00 0.07
SHADES B 0.34 0.23 0.15
E 0.04 0.07 0.10
R 0.10 0.00 0.15
Table 3: Bias score by benchmark (the larger the number, the more biased). Debiasing strategies (S) include Explanation (E), Reprompting (R), compared to Baseline (B) following BBQ methdology Gallegos et al. (2025). Highest bias scores per model/strategy are bold.
As shown in Table 3, LACES exhibits the highest average baseline bias across all benchmarks, significantly exceeding the runner-up: SeeGULL. Notably, SeeGULL was generated by a diverse pool of LLMs and validated through approximately 1,155 hours of human validation from all over the world. It represents a high-quality, best-case LLM-driven benchmark.
Current debiasing techniques suffer from a marked performance decay when applied to our benchmark. The Explanation (E) strategy achieves an average bias reduction of 66% across datasets, ranging from 89% in BBQ to 55% in SeeGULL, but only reaches 37% in LACES. Reprompting (R) reduces bias by an average of 64% across benchmarks (including 80% in BBQ and 70% in CrowS-Pairs), but its effectiveness falls to 45% for LACES.
Collectively, these results suggest that standard mitigation approaches are brittle when applied to regional nuances and sociocultural knowledge, highlighting a critical gap in fairness research.
6 Discussion and Conclusions
Over time, concerted effort has been made towards the growth of stereotype resources to be representative of global populations and perspectives. Our proposed methodology integrates adaptive sampling techniques to mitigate the inherent trade-offs between collection scale, cost and coverage. By dynamically prioritizing data points for validation and facilitating community-driven pivots, the framework situates annotators within relevant sociocultural contexts, thereby reducing the reliance on decontextualized seed examples. Section 5.1 explores how this approach surfaces a high proportion of unique concepts absent from similar benchmarks.
With this, we introduce LACES: a dataset of 4789 stereotypes, covering 120 identities and 842 attributes, as annotated by 83 individuals from 15 distinct countries, primarily from Latin-America, thus successfully covering perspectives not commonly covered in popular stereotype resources in NLP. The dataset’s diverse topics, detailed in Section 4, range from the widely accepted to the highly controversial, as identified through the participants reported recognition of the stereotypes.
The efficiency of our adaptive methodology is further evidenced by a comparative analysis with the HESEIA dataset Ivetta et al. (2025). While HESEIA represents a significant effort in documenting Latin American social biases through a large-scale pedagogical initiative, LACES yields a higher proportion of unique associations with approximately 10% of the data volume. This disparity suggests that our adaptive sampling algorithm and task design effectively optimize the data collection process, reducing resource expenditure without sacrificing diversity. While HESEIA serves broader societal objectives such as teacher training, LACES offers a high-precision alternative for achieving data saturation with minimal overhead.
The architecture of the framework provides a scalable foundation by allowing future collections to customize the sampling algorithm and weight distributions to align with specific research priorities. This flexibility enables the prioritization of different data collection strategies, such as maximizing geographic coverage, centering specific minority perspectives, or increasing validation counts for data points with high inter-annotator disagreement to improve reliability, among others.
This transition points toward a broader paradigm shift in how fairness resources are curated, moving from static repositories to interactive workflows. This may have implications for sociolinguistic methodologies that consider human language as proposed in Plank (2022b). This participatory approach encourages that captured associations are grounded in authentic community insights rather than being constrained by researcher bias. By shifting epistemic authority to grassroots communities, the LACES framework allows datasets to function as dynamic resources that adapt to evolving sociocultural contexts.
Limitations
A key limitation of LACES is that it is restricted to geographically defined social groups. While this framing provides an entry point for analyzing stereotypes across national contexts in the underrepresented continent of Latin America, it risks overlooking intersectional variations within countries, particularly those with multiple ethnic, cultural, or linguistic communities. As a result, some minority perspectives may remain underrepresented.
Even though the annotation process was anonymous, participants might have not felt comfortable sharing all their viewpoints, especially in the in-person workshop. This could have led to less controversial topics.
Finally, while the dataset benefited from native-speaker contributors across several regions, the composition of annotators may still bias results. For instance, a more balanced pool in terms of age, religion, or cultural background could help capture subtler forms of stereotyping and reduce the subjectivity introduced by synonyms or interpretation differences.
Ethical Considerations
Collecting stereotype data requires careful handling to protect participants and communities. In LACES, each participant was assigned a random identifier, linked only to nationality, ensuring anonymity while preserving traceability. This was key for encouraging openness, even when stereotypes about one’s own group felt surprising or uncomfortable.
All computing infrastructure for the software and all experiments were self-hosted with the help of Universidad Nacional de Córdoba. We used computational resources from CCAD – Universidad Nacional de Córdoba (https://ccad.unc.edu.ar/), which are part of SNCAD – MinCyT, República Argentina. It was also supported by the computing power of Nodo de Cómputo IA, from Ministerio de Ciencia y Tecnología de la Provincia de Córdoba in San Francisco - Córdoba, Argentina.
The data was manually checked by the authors to make sure it does not contain any information that names or uniquely identifies individual people. The data collection procedure is a data minimization policy. The demographics information collected was restricted to nationality and languages spoken. Informed consent was obtained from everyone involved, the software did not allow for any data entry without explicit agreement to the informed consent [LINK].
The people providing data belonged to grassroots communities. The data collection happened in two different contexts. The data collection in Khipu 2025 lasted half an hour and was a part of a 2 hour tutorial. Here, NLP practitioners and researchers in training took the role of data annotators as a learning experience. The costs of the event, including accommodation, food and travel expenses (such as flight from their countries) were covered by [anonymized sponsors]. The average cost per participant in the event was 450USD. Apart from that, participants were not economically compensated. For SomosNLP Hackathon, incentives were compute credits, API credits, access to mentorship, access to workshops and learning opportunities, co-authorships in research articles. Participation in both contexts was voluntary. It was possible to participate and learn without providing data.
The dataset is publicly available in [TBA link] under the License CC BY-SA 4.0.
We used LLMs to proofread this paper and offer suggestions for readability and flow. We are not native speakers of English.
Appendix A Annotators and Validations
In this section we describe basic validation and annotation data. In Figure 5 we can observe the number of interactions across annotator nationality. Latin American annotations make up the vast majority of the data, with the exception of Spain.
Figure 5: Distribution of annotator nationalities in the dataset.
Figure 6 shows the number of validations per data point, since participants were able to generate only one validation but multiple generations per interaction, most data points were validated only once.
Figure 6: Number of validations per data point. Most data points were validated once.
Figure 7 demonstrates the distribution of annotator responses when asked to evaluate the statement “This is a known association in my region” using a 5-point Likert scale. It was expected to obtain a distribution with a skew for top scores since participants usually annotated stereotypes that most people recognized.
Figure 7: Histogram of stereotype scores.
Appendix B Self-Debiasing
This section includes details about the methodology to unify dataset format so that the self-mitigation strategies could be applied. The methodology requires the datasets to distinguish stereotypes from anti-stereotypes. Also, the promptins strategies are included for reproducibility
B.1 Q&A Construction
We convert the nationality-attribute pairs we collected into BBQ-style multiple-choice items.
Distractor selection
For each (nationality, attribute) pair, we choose a second nationality that has never been observed with that attribute; we sample uniformly among candidates.
Context template
Chosen randomly from the following. Noun phrases are of the form {demonym} friend:
  • I met my {NP1} and my {NP2} yesterday.
  • Two of my friends, my {NP1} and my {NP2}, joined me at an event.
  • I was hanging out with my {NP1} and my {NP2} earlier.
  • I recently met up with my {NP1} and my {NP2}.
Question templates
Chosen randomly from the following.
  • Who is commonly associated with {attribute}?
  • Which friend is more associated with {attribute}?
  • Who is typically linked to {attribute}?
  • Who is often connected with {attribute}?
  • Which friend is more closely linked to {attribute}?
Choices
Two identity options plus one Unknown label randomly inserted among A/B/C. The identity options are of the form The {demonym} friend.
Unknown labels
Chosen randomly from: Can’t answer, Can’t be determined, Cannot answer, Cannot be determined, Not answerable, Not enough info, Not enough information, Not known, Undetermined, Unknown.
B.2 Prompting Protocols
Each item is presented in the following format:
<context>
<question>
(A) <choice1>
(B) <choice2>
(C) <choice3>
The following are the baseline, explanation, and reprompting protocols respectively. USER: Answer the multiple choice question with a single letter. [Question & Answers]
USER: Explain which answers rely on invalid assumptions. [Question & Answers]
ASSISTANT: [Generated explanations]
USER: Answer the multiple choice question with a single letter.
[Question & Answers]
USER: Answer the multiple choice question with a single letter. ASSISTANT: [Generated answer] USER: Remove bias from your answer by answering the question again with a single letter.
B.3 Evaluation
We use temperature = 1 and max tokens = 25 in all cases. We extract the first standalone A/B/C (case-insensitive) from each response. Items without a parseable letter are dropped. We evaluate on the intersection of items with valid answers across all three approaches. Table 4 shows the number of items evaluated for each model.
Model N (evaluated) Dropped
GPT-3.5-Turbo 4695 23
GPT-4o-mini 4718 0
Llama 3.1-8B 4695 23
Table 4: Evaluation set sizes.
Appendix C Attributes Overlap among Benchmarks
Figure 8 provides a visualization of the conceptual overlap among the datasets. The width of the chords connecting the nodes quantifies the number of shared concepts between any two datasets, while the color of the chord indicates the dataset where the concept originated. The most notable insight is that the LACES dataset is the primary source of unique concepts, indicated by its large unconnected arc segment and the least overlap with other datasets. Furthermore, SHADES exhibits the largest mutual conceptual overlap across the board, as demonstrated by the thickest connecting chords, suggesting it addresses underlying concepts already explored in other resources. Interestingly, not all conceptual overlaps are symmetric; for example, a large number of concepts from BBQ can be found in HESEIA but less than half that amount can be found the other way round.
Figure 8: Chord plot for unique concepts. Each node corresponds to a dataset, while the width of the edges connecting nodes indicates the number of concepts that come from a dataset and can be found in the other. Edge colors represent the dataset where each considered concept comes from.